OpenStudy (anonymous):

find the slope, if it exists, of the line containing the pair of points. (-14-8) and (-20-18) the slope is ____ the slope is undefined?

OpenStudy (anonymous):

18 - -8 = 26 -20 - - 14 = 6 26/6 m = 13/3

OpenStudy (anonymous):

so the answer is m = 13 over 3?

OpenStudy (anonymous):

or is 13 over 3 correct?

OpenStudy (anonymous):

13/3

OpenStudy (anonymous):

thank you